# Physics-Informed Kolmogorov-Arnold Networks for multi-material elasticity problems in electronic packaging

### v1

@@ Abstract
This paper proposes a Physics-Informed Kolmogorov-Arnold Network (PIKAN) method for analyzing elasticity problems in electronic packaging multi-material structures. The core innovation lies in replacing Multi-Layer Perceptrons (MLPs) with Kolmogorov-Arnold Networks (KANs) within the energy-based Physics-Informed Neural Networks (PINNs) framework. The method constructs admissible displacement fields that automatically satisfy essential boundary conditions and employs various numerical integration schemes to compute loss functions for network optimization. Unlike traditional PINNs that require domain decomposition and penalty terms for multi-material problems, KANs’ trainable B-spline activation functions provide inherent piecewise function characteristics that naturally accommodate material property discontinuities. Consequently, this approach requires only a single KAN to achieve accurate approximation across the entire computational domain without subdomain partitioning and interface continuity constraints. Numerical validation demonstrates PIKAN’s accuracy and robustness for multi-material elasticity problems. The method maintains high accuracy while significantly reducing computational complexity compared to domain decomposition approaches. Results confirm PIKAN’s unique advantages in solving multi-material problems and its significant potential for electronic packaging structure analysis. Source codes are available at https://github.com/yanpeng-gong/PIKAN-MultiMaterial.
@@
PINNs , Kolmogorov-Arnold networks , Multi-material structures , Electronic packaging structures , Deep energy method
@@ 1 Introduction
Moore’s Law, proposed by Gordon Moore in 1965, accurately predicted the semiconductor industry’s development trajectory and continues to drive innovation in electronic packaging technology. However, with the emergence of concepts such as ‘More than Moore’ and trends toward high-density integration and functional diversification of electronic devices, the semiconductor industry has entered a new phase highly dependent on packaging technology development [1]. Electronic packaging provides mechanical support, power distribution, electrical interconnection, and signal transmission to ensure product reliability [2]. Packaging structures typically involve multiple materials with different properties, featuring geometric discontinuities at interfaces and exhibiting multi-material, multi-scale characteristics. The reliability of such structures has become a critical design consideration, creating an urgent need for accurate and effective modeling methods in packaging structure analysis.
Numerical simulation technology has become the mainstream method for electronic packaging reliability analysis due to its advantages of easy implementation, high accuracy, and freedom from environmental constraints. In this research field, researchers have developed various numerical simulation methods to address different reliability problems in electronic packaging structures, including the finite element method (FEM) [3, 4], boundary element method (BEM) [5], FEM-BEM coupling method [6, 7], and finite difference method (FDM) [8]. Liu et al. [9] proposed a layered medium integral equation solver based on parallel computing and fast Fourier transform, achieving efficient extraction of electronic packaging network parameters. Gong et al. [10] developed an isogeometric boundary element method that accurately analyzes heat transfer problems in multi-scale electronic packaging structures with arbitrary heat sources using fewer degrees of freedom. Feng et al. [11] addressed the over-stiffness problem of traditional FEM in IGBT module multi-physics analysis through an improved NS-FEM method, enhancing computational performance. Qin and collaborators [12] used a coupling method combining finite elements and boundary elements to study elastoplastic behavior in multi-scale electronic packaging structures. Additionally, thermomechanical coupling phase field models have been applied to study damage evolution and failure mechanisms in various electronic packaging interconnect structures [13, 14].
The development of artificial intelligence (AI) technology and enhanced computational resources have provided opportunities to develop deep learning as a new approach for solving electronic packaging reliability problems. Kang et al. [15] proposed a machine learning-based thermomechanical analysis model for MMIC packaging that rapidly evaluates the impact of design parameters—including material properties, geometric features, and thermal boundary conditions-on temperature and stress distributions, providing accurate simulation alternatives for electronic packaging design. Zaghari et al. [16] investigated the collaborative design of through-hole interconnect structures in silicon and glass interposers using machine learning-based multi-objective optimization methods, analyzing the effects of different geometric parameters on electrical, thermal, and mechanical performance. Yu et al. [17] developed a GUI-based AI platform that utilizes multiple deep learning models to achieve high-precision prediction of warpage behavior in fan-out wafer-level packaging. Hou et al. [18] established a hybrid machine learning framework based on SVM and SVR for predicting bonding wire loop profiles. This approach achieves significant computational efficiency gains compared to finite element simulations, providing an effective tool for wire geometry optimization in high-density chip packaging designs. The aforementioned methods typically train neural networks based on real experimental data or high-confidence numerical calculation results (such as reliable solutions obtained through FEM) to construct surrogate models. However, training such models often requires substantial data support. In most practical engineering problems, available observational or measurement data is typically limited, which may cause trained models’ prediction results to deviate from known physical laws or lose physical consistency [19], thereby reducing prediction accuracy and constraining model generalization capability and interpretability. Machine learning techniques, when coupled with conventional computational methods, offer tremendous potential to revolutionize modeling and simulation in engineering and the sciences, opening new avenues for next-generation computer-aided analysis and design [20].
@@
To reduce neural networks’ dependence on high-quality data, researchers embed physical laws into loss functions to enhance models through physics-informed constraints. Physics-Informed Neural Networks (PINNs), proposed by Raissi et al. [21] in 2019, became the pioneering work in this direction. The implementation of PINNs benefits from the mature application of automatic differentiation functionality in machine learning frameworks such as PyTorch and TensorFlow. As a meshfree method using neural networks as approximation functions, PINNs inherit the advantages of meshfree methods in handling mesh distortion and large deformation problems [22] and can solve PDE problems that traditional numerical algorithms can address. PINNs encompass two forms: strong-form original PINNs and energy-form PINNs. Samaniego et al. [23] introduced the Deep Ritz Method proposed by E and Yu [24] into computational mechanics, forming energy-form PINNs called the Deep Energy Method (DEM). Strong-form PINNs construct loss functions by directly combining PDEs through weighted residual methods, while DEM constructs loss functions based on variational principles such as the minimum potential energy principle in mechanics. In electronic packaging structure analysis, although there are limited studies using traditional data-driven surrogate models based on physical simulation data [25], research directly applying PINNs to such simulations has rarely been reported.
Multi-Layer Perceptrons (MLPs) serve as the core architecture of PINNs and the standard model for approximating nonlinear functions in machine learning. The universal approximation theorem [26] demonstrates that MLPs can approximate any continuous function with arbitrary precision given sufficient hidden neurons and appropriate weight configurations. However, their limited parameter efficiency significantly increases computational complexity and causes convergence difficulties when handling complex problems [27]. Recently, Liu et al. [28] proposed Kolmogorov-Arnold Networks (KANs) as a novel alternative to MLPs. While structurally similar to MLPs, KANs use learnable one-dimensional functions [29], to replace linear weight parameters, with nodes simplified to perform input summation. KANs integrate spline function approximation capabilities with MLP network architecture, typically achieving model expression with smaller computational graphs. Given that MLPs suffer from large parameter scales and limited interpretability, while KANs offer fewer parameters and B-spline-based activation functions that align more naturally with numerical PDE solving algorithms, integrating KANs into PINNs to replace MLPs is theoretically well-motivated.
@@
Electronic packaging structures are typically composed of heterogeneous materials with different properties, exhibiting multi-material, multi-interface structural characteristics. When applying PINNs to solve multi-material problems, the current mainstream approach uses different neural networks to describe physical fields in different material domains. Diao et al. [30] proposed a strong-form PINNs framework for solving multi-material solid mechanics problems by partitioning the computational domain through domain decomposition, utilizing independent sub-networks to characterize mechanical behavior of different materials, and introducing interface regularization terms to handle contact relationships and discontinuities. Wang et al. [31] focused on variational problems and proposed a conservative energy method based on subdomain neural networks for solving heterogeneous material and complex geometry problems. However, PINNs based on subdomain partitioning have inherent limitations. To accurately describe interface contact relationships, additional hyperparameters must be introduced, which typically rely on empirical parameter tuning. Although research has explored hyperparameter optimization strategies [32, 33], obtaining optimal parameter combinations remains challenging when solving specific problems. Furthermore, this method requires independent optimization of neural networks for each subdomain, which significantly increases training complexity and often causes the optimization process to converge to local rather than global optima.
Based on the aforementioned research background, this paper adopts a Physics-Informed Kolmogorov-Arnold Network (PIKAN) method for solving multi-material problems in solid mechanics and extends its application to electronic packaging multi-material structure analysis. This method replaces traditional multi-layer perceptrons with KANs and combines them with the DEM. Compared to strong-form PINNs, DEM has stronger physical interpretability and avoids high-order differential operations, resulting in higher computational efficiency and superior accuracy [34]. In PIKAN, admissible displacement fields are constructed [31, 32] to strictly satisfy essential displacement boundary conditions of geometric structures. Furthermore, benefiting from KANs’ trainable one-dimensional activation functions and inherent piecewise function characteristics, this method requires no subdomain partitioning during implementation and can approximate the entire solution domain using only a single KAN. This characteristic effectively circumvents the need for additional interface information and excessive hyperparameters, significantly reducing training difficulty and computational cost. Numerical example verification demonstrates that the proposed PIKAN method performs well when handling structures containing multiple materials and different interface configurations, broadening the application of PINNs in multi-material problems. Meanwhile, its application in electronic packaging multi-material structure analysis validates the method’s potential and value in electronic packaging structure reliability analysis.
@@ 2 Deep energy method-based computational solid mechanics
This section reviews the fundamental equations of elasticity problems, then introduces PINNs with energy-based loss functions for solving such problems. Finally, we present the construction of admissible displacement fields that naturally satisfy essential displacement boundary conditions.
2.1 Fundamental equations of elasticity problems
The governing equation at any point $\mathbf{x}$ within the elastic body domain $\Omega$ is
$\sigma_{ij,j}+\bar{f}_{i}=0,\quad\mathbf{x}\in\Omega$ (1)
where $\sigma$ is the Cauchy stress tensor, $\bar{f}_{i}$ is the body force per unit volume, and $(\cdot)_{,j}=\partial(\cdot)/\partial x_{j}$. For isotropic linear elastic materials, the stress-strain relationship is expressed as
$\sigma_{ij}=\lambda\delta_{ij}\varepsilon_{kk}+2\mu\varepsilon_{ij}=D_{ijkl}\varepsilon_{kl}$ (2)
where $\delta_{ij}$ is the Kronecker delta, $\varepsilon_{ij}$ is the strain tensor, $D_{ijkl}$ is the elasticity tensor, and $\lambda$ and $\mu$ are the Lamé constants characterizing material properties.
Under the small deformation assumption, the strain-displacement relationship is expressed as
$\varepsilon_{ij}=\frac{1}{2}(u_{i,j}+u_{j,i})$ (3)
where $u_{i}$ represents the $i$-th displacement component. The governing equations are closed through boundary conditions. The complete boundary of the elastic body $\Omega$ is defined as $\Gamma$. A portion of the boundary where displacements are prescribed is called the essential (Dirichlet) boundary, denoted as $\Gamma_{u}$. The portion where tractions are prescribed is called the natural (Neumann) boundary, denoted as $\Gamma_{\sigma}$. These two boundary segments constitute the complete boundary
$\Gamma_{u}\cup\Gamma_{\sigma}=\Gamma,\quad\Gamma_{u}\cap\Gamma_{\sigma}=\emptyset$ (4)
The displacement boundary conditions are expressed as
$u_{i}=\bar{u}_{i},\quad\mathbf{x}\in\Gamma_{u}$ (5)
where $\bar{u}_{i}$ is the prescribed displacement on the corresponding boundary. The traction boundary conditions are expressed as
$p_{i}=\sigma_{ij}n_{j}=\bar{p}_{i},\quad\mathbf{x}\in\Gamma_{\sigma}$ (6)
where $n_{j}$ is the $j$-th component of the outward unit normal vector, and $\bar{p}_{i}$ is the prescribed traction on the corresponding boundary.
@@ 2.2 Deep energy method for solving elasticity problems
For elastic systems, DEM utilizes the principle of minimum potential energy from variational elasticity theory. Based on the principle of minimum potential energy, an energy-based loss function that utilizes the total potential energy of the system as the optimization objective has been proposed within the variational PINNs framework [23]. The core idea is to directly construct the total potential energy of the elastic body as the loss function for neural network training. For problems without body forces, the specific mathematical expression of this loss function is
$\mathcal{L}=\Psi_{\text{in}}-\Psi_{\text{ex}}$ (7)
where $\Psi_{\text{in}}$ and $\Psi_{\text{ex}}$ represent the internal strain energy and the potential energy of external forces, respectively, with their specific expressions given by
$\Psi_{\text{in}}=\int_{\Omega}\frac{1}{2}D_{ijkl}\varepsilon_{ij}\varepsilon_{kl}\mathrm{d}\Omega=\int_{\Omega}\frac{1}{2}\sigma_{ij}\varepsilon_{ij}\mathrm{d}\Omega$ (8)
$\Psi_{\text{in}}=\int_{\Gamma_{\sigma}}\bar{p}_{i}u_{i}\mathrm{d}\Gamma$ (9)
Therefore, the energy-based loss function in Eq. (7) can be written as
$\mathcal{L}=\int_{\Omega}\frac{1}{2}\sigma_{ij}\varepsilon_{ij}\mathrm{d}\Omega-\int_{\Gamma_{\sigma}}\bar{p}_{i}u_{i}\mathrm{d}\Gamma$ (10)
While DEM shares similarities with FEM in using energy-based variational weak forms, the key distinction lies in its direct optimization of the total potential energy functional, eliminating the need for preliminary derivation and discretization of weak form governing equations required in FEM.
In DEM, the displacement field $\mathbf{u}(\mathbf{x})$ satisfying the prescribed essential boundary conditions is approximated through a neural network
$\mathbf{u}(\mathbf{x})\approx\mathbf{F}(\mathbf{x};\boldsymbol{\theta})$ (11)
$\text{s.t. }\mathbf{u}(\mathbf{x})=\overline{\mathbf{u}}(\mathbf{x}),\quad\mathbf{x}\in\Gamma_{u}$
where $\mathbf{F}(\cdot)$ represents the neural network mapping, $\overline{\mathbf{u}}(\mathbf{x})$ denotes the essential boundary conditions, and $\boldsymbol{\theta}$ are the trainable parameters of the neural network.
Based on the principle of minimum potential energy, among all kinematically admissible displacement fields $\mathbf{u}(\mathbf{x})$ that satisfy essential boundary conditions, we seek the solution that minimizes the total potential energy functional
@@
where $\psi(\boldsymbol{\varepsilon}(\mathbf{u}(\mathbf{x})))$ is the strain energy density function, which is a function of the strain tensor $\boldsymbol{\varepsilon}(\mathbf{u}(\mathbf{x}))$.

### v2

@@ Abstract
This paper proposes a Physics-Informed Kolmogorov-Arnold Network for analyzing elasticity problems in multi-material electronic packaging structures. The method replaces traditional Multi-Layer Perceptrons with Kolmogorov-Arnold Networks within an energy-based Physics-Informed Neural Network framework. By constructing admissible displacement fields satisfying essential boundary conditions and optimizing network parameters through numerical integration, the proposed method effectively handles material property discontinuities. Unlike traditional methods that require domain decomposition and interface constraints for multi-material problems, Kolmogorov-Arnold Networks’ trainable B-spline activation functions provide inherent piecewise characteristics. This capability stems from B-splines’ local support, which enables effective approximation of discontinuities despite their individual smoothness. Consequently, this approach enables accurate approximation across the entire domain using a single network and simplifying the computational framework. Numerical experiments demonstrate that the proposed method achieves excellent accuracy and robustness in multi-material elasticity problems, validating its practical potential for electronic packaging analysis. Source codes are available at https://github.com/yanpeng-gong/PIKAN-MultiMaterial.
@@
Physics-Informed Neural Network , Kolmogorov-Arnold networks , Multi-material structures , Electronic packaging structures , Deep energy method
@@ 1 Introduction
Electronic packaging provides critical mechanical support, electrical interconnection, and thermal management for high-density integrated electronic devices Moreau [2012]. Its reliability analysis involves challenges including multi-material systems, multi-scale phenomena, and interface discontinuities. Numerical simulation techniques, benefiting from easy implementation, high accuracy, and environment-independent operation, have become mainstream methods for electronic packaging reliability analysis, including finite element method (FEM) Kim et al. [2020], Feng et al. [2023], boundary element method (BEM) Gong et al. [2022], Yu et al. [2021], coupled algorithms Gong et al. [2025a], Qin et al. [2022], finite difference method (FDM) Zhang [2004], and phase-field models Gong et al. [2024, 2025b], among others, with each method advancing through specific applications. However, facing the growing design complexity and computational efficiency demands of modern packaging structures, traditional methods still have inherent limitations in computational cost, local field resolution, and physical modeling, creating urgent demand for more efficient and flexible computational frameworks.
The development of artificial intelligence (AI) technology and enhanced computational resources have provided opportunities for deep learning as a new approach to electronic packaging reliability problems. Researchers have successfully applied machine learning to construct surrogate models for rapid thermal-mechanical coupling behavior assessment Kang et al. [2023], geometric parameter co-design Zaghari et al. [2025], and process prediction Yu et al. [2025], demonstrating significant computational efficiency advantages. However, these data-driven methods typically require substantial high-confidence training data; in scenarios with limited engineering data, their predictions may deviate from physical laws, affecting model generalization and physical consistency Yuan et al. [2022]. Therefore, combining machine learning with traditional computational methods offers tremendous potential for transforming modeling and simulation in engineering and scientific domains, opening new pathways for next-generation computer-aided analysis and design Rabczuk and Bathe [2023].
@@
In recent years, deep learning methods have provided new numerical paradigms for solving solid mechanics problems by avoiding traditional mesh generation and improving computational efficiency Linghu et al. [2025a], Heidenreich et al. [2024]. To reduce neural networks’ dependence on high-quality data, researchers embed physical laws into loss functions to enhance models through physics-informed constraints. Physics-Informed Neural Networks (PINNs), proposed by Raissi et al. Raissi et al. [2019] in 2019, became the pioneering work in this direction. The implementation of PINNs benefits from the mature application of automatic differentiation functionality in machine learning frameworks such as PyTorch and TensorFlow. Compared to traditional numerical algorithms, PINNs leverage the powerful approximation capabilities of neural networks to provide a new paradigm for solving complex problems that conventional numerical methods struggle to address Cuomo et al. [2022]. Furthermore, PINNs demonstrate significant potential in handling high-dimensional problems, as their loss functions are typically constructed through Monte Carlo sampling, which is inherently an effective tool for high-dimensional integration. PINNs encompass two forms: strong-form original PINNs and energy-form PINNs. Samaniego et al. Samaniego et al. [2020] introduced the Deep Ritz Method proposed by E and Yu E and Yu [2018] into computational mechanics, forming energy-form PINNs called the Deep Energy Method (DEM). Strong-form PINNs construct loss functions by directly combining PDEs through weighted residual methods, while DEM constructs loss functions based on variational principles such as the minimum potential energy principle in mechanics. The flexibility and potential of the PINN framework continue to drive its development and application in various cutting-edge engineering fields Linghu et al. [2025a], Ma et al. [2025], Li et al. [2025], Chen et al. [2025], Niaki et al. [2021], Xiong and Zhao [2025], Linghu et al. [2025c]. In electronic packaging structure analysis, although there are limited studies using traditional data-driven surrogate models based on physical simulation data Yao et al. [2022], research directly applying PINNs to such simulations has rarely been reported.
Within the standard PINN framework, multi-layer perceptrons have been widely adopted due to their structural simplicity and universal approximation properties Hornik et al. [1989]. However, the framework is inherently flexible, and the approximation function is not limited to MLPs. Research has successfully extended to various architectures such as RBFs Gao et al. [2025], Wavelets Zhu et al. [2026], and others. Recently, Kolmogorov-Arnold Networks (KANs), proposed by Liu et al. Liu et al. [2024], employ trainable one-dimensional activation functions that simultaneously replace linear weights and provide nonlinear activation behavior, with nodes simplified to perform only input summation, have also been introduced into the PINN framework, rapidly spawning multiple improved studies focusing on fair comparisons with MLPs Shukla et al. [2024], architectural innovations Wang et al. [2025b], Jacob et al. [2025], and other aspects. When dealing with PDE problems involving high-dimensional or complex solution spaces, MLPs often require extremely wide or deep networks to achieve the required accuracy, leading to an enormous number of parameters. This increases optimization difficulty and computational cost Han et al. [2018], while KANs provide an alternative parameterization approach through their spline-based learnable activation functions. Meanwhile, the spectral bias problem in MLP training may lead to insufficient learning capability for high-frequency components of PDE solutions, whereas the flexible function representation provided by KANs has been indicated to be more conducive to capturing multi-scale features of solutions and alleviating spectral bias phenomena Wang et al. [2025a], Meshir et al. [2025]. This provides a new potential avenue for solving PDE problems with multi-scale characteristics or solutions with sharp transitions. It should be noted that the completeness of the universal approximation theory for classical KANs is still being refined, with only a few variants rigorously proven to date Toscano et al. [2025b]. Nevertheless, the excellent computational performance demonstrated by this architecture in a wide range of scientific computing problems motivates us to further explore its application potential in PDE solving.
@@
Electronic packaging structures are typically composed of heterogeneous materials with different properties, exhibiting multi-material, multi-interface structural characteristics. For multi-material structures with sharp interfaces, the meshfree nature of PINNs enables them to avoid the mesh distortion issues that may arise in FEM. When applying PINNs to solve multi-material problems, the current mainstream approach uses different neural networks to describe physical fields in different material domains. Diao et al. Diao et al. [2023] proposed a strong-form PINNs framework for solving multi-material solid mechanics problems by partitioning the computational domain through domain decomposition, utilizing independent sub-networks to characterize mechanical behavior of different materials, and introducing interface regularization terms to handle contact relationships and discontinuities. Wang et al. Wang et al. [2022b] focused on variational problems and proposed a conservative energy method based on subdomain neural networks for solving heterogeneous material and complex geometry problems. However, PINNs based on subdomain partitioning have inherent limitations. To accurately describe interface contact relationships, additional hyperparameters must be introduced, which typically rely on empirical parameter tuning. Although research has explored hyperparameter optimization strategies Sheng and Yang [2021], Wang et al. [2022a], obtaining optimal parameter combinations remains challenging when solving specific problems. Furthermore, the domain decomposition method requires independent optimization of neural networks for each subdomain, which increases training complexity and often causes the optimization process to converge to local rather than global optima.
Although Physics-Informed Kolmogorov-Arnold Networks (PIKANs) have received increasing attention Liu et al. [2024], Shukla et al. [2024], Wang et al. [2025b], Jacob et al. [2025], Toscano et al. [2025a], Mostajeran and Faroughi [2025], Faroughi and Mostajeran [2025], the distinctive contribution of this study lies in developing a straightforward PIKAN solution framework specifically tailored for multi-material solid mechanics problems with various interface configurations commonly encountered in electronic packaging structures, integrating KANs with the DEM to replace traditional MLPs. To our knowledge, we present and validate for the first time this integrated approach for electronic packaging multi-material structure analysis, focusing on the persistent challenge of modeling multi-material structures with inherent interface discontinuities. Within the PINN framework, the strong form has broader generality, but its application typically requires extensive hyperparameter tuning. In contrast, the Deep Energy Method, based on variational principles, transforms the governing equations into an extremum problem of energy functionals. This not only reduces the number of required hyperparameters, but also typically achieves higher computational accuracy and efficiency due to the reduced order of functional derivatives Li et al. [2021], though its generality is usually weaker than the strong form. For elastic analysis problems of multi-material structures, this study aims to develop a more concise solution paradigm, and therefore adopts the DEM with fewer hyperparameters. It should be noted that theoretical analysis of numerical convergence and stability for PINNs and their variational (energy-based) forms has made certain progress Xu et al. [2024], Gazoulis et al. [2025]. For instance, rigorous convergence proofs have been established for specific higher-order multi-scale PINN methods in composite mechanics Linghu et al. [2025b]. However, for the emerging KAN architecture and its applications in physics-informed learning, rigorous convergence theory for KAN/PIKAN architectures remains an active research frontier.
Based on the above research background, this paper employs the PIKAN method to solve multi-material problems in solid mechanics and extends its application to electronic packaging multi-material structure analysis. This work constructing admissible displacement fields Wang et al. [2022b], Sheng and Yang [2021] that strictly satisfy essential boundary conditions. The main contributions of our work include:
  • 1.
    Development of PIKAN methodology that replaces traditional MLPs with KANs in deep energy method, leveraging KANs’ inherent piecewise function characteristics to naturally accommodate material property discontinuities.
  • 2.
    Elimination of domain decomposition requirements by using a single KAN to approximate displacement fields across the entire computational domain without subdomain partitioning or interface continuity constraints.
  • 3.
    Comprehensive validation through diverse examples including cantilever beams with different interface geometries and other electronic packaging models, demonstrating superior accuracy compared to traditional domain decomposition approaches.
  • 4.
    Systematic investigation of numerical integration schemes and hyperparameter sensitivity analysis, with complete algorithm implementation details and source code to be provided in our GitHub repository, facilitating practical application, comparison, and further development of PIKANs for multi-material problems.
@@ 2 Deep energy method-based computational solid mechanics
This section introduces the Deep Energy Method (DEM) for solving elasticity problems. Then we present the construction of admissible displacement fields that naturally satisfy essential displacement boundary conditions.
@@ 2.1 Deep energy method for solving elasticity problems
For elastic systems, based on the principle of minimum potential energy in variational elasticity theory, the Deep Energy Method (DEM) constructs the total potential energy functional of the system as a loss function for optimization Samaniego et al. [2020]. For problems without body forces, the specific mathematical expression of this loss function is
$\mathcal{L}=\Psi_{\text{in}}-\Psi_{\text{ex}}=\int_{\Omega}\frac{1}{2}\sigma_{ij}\varepsilon_{ij}\mathrm{d}\Omega-\int_{\Gamma_{\sigma}}\bar{p}_{i}u_{i}\mathrm{d}\Gamma$ (1)
where $\Psi_{\text{in}}$ and $\Psi_{\text{ex}}$ represent the internal strain energy and external force potential energy, respectively; $\sigma_{ij}$ and $\varepsilon_{ij}$ denote the stress and strain tensors, respectively; $u_{i}$ represents the $i$-th displacement component; and $\bar{p}_{i}$ is the traction applied on the natural boundary. DEM shares similarities with FEM in adopting energy-based variational weak forms, but its key distinction lies in the fact that DEM directly optimizes the total potential energy functional, thereby eliminating the steps of preliminary derivation and discretization of the weak form of governing equations required by the FEM. For the electronic packaging multi-material structures of interest, we focus on mechanical responses dominated by external mechanical loads and material interface effects. The body force term is omitted in the energy functional construction, based on engineering simplification appropriate for the scale and loading characteristics of the studied problems Bai et al. [2023], Ye et al. [2022]. The proposed framework can straightforwardly incorporate body force terms to handle more general scenarios when needed.
Among all kinematically admissible displacement fields $\mathbf{u}(\mathbf{x})$ that satisfy the essential boundary conditions, we seek the solution that minimizes the total potential energy functional
@@
where $\mathbf{u}(\mathbf{x})$ is approximated through a neural network as $\mathbf{u}(\mathbf{x})\approx\mathbf{F}(\mathbf{x};\boldsymbol{\theta})$, $\mathbf{F}(\cdot)$ denotes the neural network mapping, and $\psi(\boldsymbol{\varepsilon}(\mathbf{u}(\mathbf{x})))$ is the strain energy density function, which is a function of the strain tensor $\boldsymbol{\varepsilon}(\mathbf{u}(\mathbf{x}))$.